OpenStudy (anonymous):

calculus exam tomorrow please help me, find the sum of the series :

OpenStudy (anonymous):

\[ \sum_{k=1}^{\infty}1/k(k+3)\]

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}& \sum_{1}^{\infty} \dfrac{1}{k(k+3)}\hspace{.33em}\\~\\ & \implies \sum_{1}^{\infty} \dfrac{3}{3k(k+3)}\hspace{.33em}\\~\\ & \implies \sum_{1}^{\infty} \dfrac{k+3-k}{3k(k+3)}\hspace{.33em}\\~\\ & \implies \sum_{1}^{\infty} \dfrac{1}{3k}-\dfrac{1}{3(k+3)}\hspace{.33em}\\~\\ & \implies \dfrac{1}{3}\sum_{1}^{\infty} \dfrac{1}{k}-\dfrac{1}{(k+3)}\hspace{.33em}\\~\\ \end{align}}\) now try to put some values up to 9 or 10 and see if it gives telescooping series

OpenStudy (mathmath333):

https://en.wikipedia.org/wiki/Telescoping_series

OpenStudy (anonymous):

@mathmath333 thank you so much..